media%2F406%2F406a012c-257a-42ea-863f-69
Answer #1

given , volume V = 2.5 litres = 2.5*10^-3 m^3 (squinch)
volume of serroboam, v = 8.1 us gallons
now 1 us gallon = 3.78541 litres
so in 1 serroboam, 8.1*3.78541 litres = 8.1*3.78541/2.5 squinches
so in 1 serroboam there are 12.264 squnches


height of monument, h = 192 m
angle of line of siognt, theta = 7.2 degree above h orizontal
let the distance from the base of the monument be d
then tan(theta) = h/d
d = h/yan(theta) = 192/tan(7.2) = 1519.836 m
d = 1.519 km